Mathematics 133 – Intensive Calculus for Sciences General Information on Maple in HA 408

Launching Maple

To get into Maple, you will need to follow these directions.

1) If your computer is already turned on and the Windows desktop is visible, go to step 2 below. Otherwise, turn the computer and monitor on, and wait a few minutes while the computer “boots up.” If you do this, eventually you will see a prompt to log onto the Novell campus network. Enter your campus network username and password, and go to step 3.

2) If your computer is already turned on and the Windows desktop is visible, but you have not yet logged on, click the Start button. In the Start menu you will see either “Log off xxx”, where “xxx” is the username of whoever was on before, or else just “Log off...”. In the first case, it means that someone else left in the middle of a session. You do not want to do that. It means that your P: drive, your email, etc. are readable by whoever sits down at that computer next. Always log off when you are finished in a session. If someone else is logged on, log them off. Then click the Start button again. Now that item in the Start menu should just say “Log off...”. If you select that, you will get the Novell login dialog to log on yourself. Enter your username and password.

3) You should now see the Windows desktop. Double click on Maple 7 shortcut icon.

4) After a few more seconds you should see a new Maple 7 window, with a “subwindow” marked Untitled (1) opened inside it. The Untitled (1) window is a blank Maple worksheet.

Maple Worksheets

Worksheets are integrated mathematical documents where any or all of the following can be done:

1) you can type in commands from the keyboard at the input prompts [>, to ask Maple to perform many different kinds of calculations,

2) Maple will generate output (numerical values, symbolic formulas, and graphics),

3) you can modify commands, generating new output, in our session today.

4) you can enter text to annotate and explain the results of computations. Take a few seconds and notice the features of this window – especially the “tool bar” across the top with the icons for various operations, the “scroll bar” on the right that you can use to move around within the worksheet, to see previous input and output lines, etc.

Input and Text Regions in a Worksheet

A new input prompt will be generated automatically at the bottom of a worksheet each time you enter a command and execute it at the end of a worksheet in progress. You can also insert a command prompt and an input region at any point in the worksheet by placing the cursor at the desired location and pressing the toolbar button marked with [>. (After executing an inserted command like this, Maple will drop to the next input line, so to insert several input lines in the middle of a worksheet, you will need repeat the above.)

As we mentioned before, Maple worksheets can contain text as well as commands and output. To create a text region you can press the [> button above, then the button marked by a capital T (this changes the region into a text region). Any text can now be entered, and this will be treated as an “inert” comment. That is, it will appear as you enter it when you print out the worksheet, and it will not be treated as Maple input.

After you create a text region, if you press the capital sigma (Σ) toolbar button with the cursor placed on that region, you can type in a Maple expression in input format, and have it displayed (in “inert” form) in usual mathematical notation (with raised exponents, no asterisks for multiplication, etc. This is sometimes useful for creating “pretty” worksheets!

Saving and Reloading your Maple Worksheets

When you begin working on a worksheet, you will usually want to save your work every once in a while in case a computer problem develops, or in case you need more than one lab session to complete the work you are doing. This can be done most directly by saving to your network P: drive. (You can also save to a 3 1/2 inch High Density IBM format diskette.) Follow these directions:

1) If you are using a diskette, insert it in the diskette drive (it’s called the a: drive on all the lab machines). Otherwise go directly to step 2.

2) In Maple, select the SAVE option from the FILE pull-down menu.

3) If you are saving your work for the first time, you will see a SAVE AS dialog box. In the Drives box, highlight the A: or P: drive as appropriate, then go to the File Name box, and type in a name for the file where the worksheet will be stored (any string of letters and digits – no spaces – no more than 8 characters long is OK) type in a period after the file name you chose and then the “extension” mws for Maple work sheet. For instance, a good choice might be something like labday1.mws if you are working on the first lab assignment. Then click the OK button to write the worksheet to the P: drive or to your diskette. (You will only need to type the filename in once in a session. Subsequent saves just update the file.)

4) When you have the worksheet saved as you want it, you can remove the diskette and exit Maple.

5) To update the worksheet further in a later session, you can read the worksheet back into Maple using the OPEN option from the FILE Menu. Get into Maple as above, (if the worksheet is stored on diskette, insert it in the a: drive), and select OPEN from the FILE Menu. Maple will prompt you as above for the name of the worksheet with a dialog box very like the SAVE AS box described above. Highlight the A: or P: drive as appropriate, then the worksheet you want, and click OK.

Important note: If you read in a previously saved worksheet, the output from the last session will be displayed, but any values saved in variables in that session will not be known this time. You will need to reexecute any such command lines in this session. This can be done by pressing RETURN on each input line again, or by selecting the EXECUTE WORKSHEET option from the EDIT toolbar menu.

Printing in HA 408

To print a Maple worksheet in HA 408, you will need to log onto the campus network before launching Windows and Maple. This is done by typing in your network username and password at the prompts that come up automatically when you boot up. Then get into Windows and Maple as above. OPEN the worksheet you want to print, and press the toolbar icon that looks like a printer, and press OK on the PRINT dialog box (the settings should be set up correctly to print automatically). Your output should appear shortly on the printer at the end of the lab. If you have problems, or if the printer runs out of paper, come and get me or another math faculty member for help. (If several print jobs are sent at the same time, you may need to wait a short time.)

Getting Out

When you leave, quit the Maple window (FILE/EXIT), and log off the campus network by clicking the Start button and selecting “Log off.”